what value makes x^2-9/x^2+7x+10 undefined?
Answer by chriswen(106)

The expression is undefined if the bottom equals 0.
x^2+7x+10=0
(x+5)(x+2)=0
x+5=0 or x+2=0
x=-5 or x=-2
If x=-5 or x=-2 then the bottom will equal 0. Then, the expression would be undefined.